Kabbed:

Your boss hands you the monthly data that show the number of orders coming in to and out of the warehouse. The data are in the table below. Explain to your boss, in complete sentences, the solution to this system and what the solution represents.

Kabbed:

http://prntscr.com/n9bqy4

Kabbed:

@Vocaloid

Kabbed:

@Pixel

Kabbed:

@TheSmartOne

Kabbed:

Please help me

TheSmartOne:

Hi there! Welcome to QuestionCove! I have a few minutes before I need to leave but I'll try to show you how to answer your question: Your month is your x-coordinate The no. of orders IN is your y-coordinate. For example (1, 3) Likewise, we will have another set of points where Your month is your x-coordinate The no. of orders OUT is your y-coordinate You will first need to calculate the lines of each set of points. Start by first finding the slope. \(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\) To find the slope, you only need 2 points. After you find the slope, plug that into y = mx + b where m is the slope and b is the y-intercept So you should have a number replacing 'm' After that, plug in one of your points (it can be one of the points you used to calculate your slope) and plug that into the equation and solve for b Once you get your two equations for both data the orders OUT and the orders IN, you set them equal to each since they both will be in the form of y=mx+b And so you can solve for x which would be the solution to the system

Kabbed:

Slope = 3 - 3/6 - 4

Kabbed:

slope = 3 - 3/1 - 1

Kabbed:

3, 3 are the order in and out and 1 is the month of Janurary

Kabbed:

@563blackghost help I don't understand

Kabbed:

Slope = I don't know what to do or what numbers to write

Kabbed:

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\)